OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (1, 9) and (-1, 11)?

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (11-9)/(-1-1) m = (2)/(-2) m = -1 ------------------------------ y - y1 = m(x - x1) y - 1 = -1(x - 9) y - 1 = -(x - 9)

jimthompson5910 (jim_thompson5910):

You can do further simplification, but then it wouldn't be in point-slope form any more